OpenStudy (anonymous):

Christa buys 14 flowerpots. Some are small and cost $2.40 each. The rest are large and cost $5.60 each. She spends a total of $49.60. Which equation models this situation? Let s represent the number of small flowerpots she buys. A. 5.6s + 2.4(s + 14) = 49.6 B. (2.4 + 9.6)(14 – s) = 49.6 C. 2.4s + 5.6(14 – s) = 49.6 D. 5.6s + 2.8(14) = 49.6

OpenStudy (mathstudent55):

If the total number of flowerpots is 14, then small + large = 14 Ok?

OpenStudy (anonymous):

i think it is D am i right

OpenStudy (mathstudent55):

We'll get there soon. Just follow along what I am doing.

OpenStudy (mathstudent55):

small pots + large pots = 14 Right?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now let's see how many large pots there are: small pots + large pots = 14 They state that you should let s = number of small pots, so we have s + large pots = 14 Subtract s from both sides: large pots = 14 - s

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now we know this: large pots = 14 = s small pots = s

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now we use the prices. If one small pot costs $2.40, how much do s small pots cots?

OpenStudy (mathstudent55):

(Hint: multiply the number of pots by the price of one pot.)

OpenStudy (anonymous):

so $2.40

OpenStudy (mathstudent55):

2.40 * 1 is the price of 1 pot 2.40 * 2 = price of 2 pots 2.40 * 3 = price of 3 pots 2.40 * s = price of s pots The price of s pots is: 2.40s

OpenStudy (anonymous):

so the answer is C?

OpenStudy (mathstudent55):

Now we find the price of the large pots. One large pot costs $5.60 There are (s - 14) large pots, so the price of all the large pots is: 5.60(s - 14)

OpenStudy (mathstudent55):

price of small pots + price of large pots = total price 2.4s + 5.6(s - 14) = 49.6 You are correct. This is choice C.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

You're welcome.